# Incremental Composition of Learned Control Barrier Functions in Unknown Environments

### v1

@@ III-A Obtaining Expert Demonstrations Online
Figure 2: Left-to-Right: Incrementally learning a global CBF for the Dubins car, while safely exploring an unknown environment under input constraints. LiDAR scans are taken (and CBFs computed) where the trajectory color changes. To showcase the correctness of our local CBFs, each trajectory is constrained by the last-learned local CBF, whose zero-level set is shaded. At each $(q^{1},q^{2})$, red arcs denote unsafe angles $h(q^{1},q^{2},\theta)<0$. Blue arcs denote $h(q^{1},q^{2},\theta)\geq 0$. Red arcs that are not on obstacles reflect the fact that unexplored states must be considered unsafe.
The Learning-CBF QP described in Section II-A provides us with a powerful framework to learn valid CBFs from demonstrations. In order to apply the framework, we need access to a dataset of expert demonstrations containing safe state-action pairs (cf. Definition 2). However, this may not be the case as we freely explore the state space. For example, suppose our measurement map is implemented via a LiDAR sensor. While this sensor provides us with a point cloud of obstacles, it does not provide us with expert control inputs that safely avoid said obstacles.
We provide two remedies for the lack of safe inputs. Our first solution is based on the dual norm. Suppose that the input set $U$ is a norm ball of the form $\|u\|\leq u_{\max}$. Then, the CBF derivative constraint $\sup_{u\in U}\langle\nabla h(x),\ f(x)+g(x)u\rangle\geq-\alpha(h(x))$ and hence the constraint Eq. 3 in the Learning-CBF QP can be rewritten in terms of the dual norm $\|u\|^{\star}=\sup_{x\in\mathbb{R}^{d_{n}}}\{\langle x,u\rangle:\|x\|\leq 1\}$ as follows [22]:
$\langle\nabla h(x),f(x)\rangle+u_{\max}\|g(x)^{T}\nabla h(x)\|^{\star}+\alpha(h(x))\geq\gamma_{\text{dyn}}.$
Observe that the above expression does not require any input demonstrations $u$. However, there are two limitations to this approach. Firstly, the resulting Learning-CBF QP is nonconvex. Secondly, sensors may not fully characterize the safe states (i.e. $d_{q}<d_{x}$), e.g., consider a car with a LiDAR sensor. The LiDAR provides safe and unsafe positions, but not the safe steering angles and velocities which are needed to avoid system states that inevitably lead to a collision.
To address the limitations of the dual norm, our second approach is based on the notion of a safety oracle. Given a safe set returned by a sensor, we define a safety oracle to be a function that takes this safe set as input and returns safe state-action pairs for which there exists a control signal that keeps the system in the aforementioned safe set for all time.
@@
Some examples of safety oracles are Hamilton-Jacobi reachability analysis [9], safe MPC [23], and obstacle-aware shooting methods [24]. Our proposed approach allows us to efficiently use local oracle computations in two ways. Firstly, while oracle computations may only provide safe state-action pairs for a finite set of states, the learned locally-valid CBF interpolates between the given datapoints (cf. Definition 2), yielding safety guarantees over the entire local domain. Secondly, oracle computatations may be costly. By synthesizing local safety computations into a composable CBF, we can store and combine them in a principled way, circumventing the need to recalculate the safety of already-certified trajectories. The following example illustrates how our approach leverages a safety oracle.
Figure 3: The global CBF for the Dubins car. Observe that the magenta trajectory of the system under the global CBF interpolates the behavior of the system under the locally learned CBFs shown in Fig. 2.
Dubins car: Consider dynamics $\dot{x}=(\dot{q}^{1},\dot{q}^{2},\dot{\theta})=(v\cos\theta,v\sin\theta,u)$, with position $q:=(q^{1},q^{2})$ and a fixed velocity $v$. The system operates in an unknown environment with access to a LiDAR sensor that returns a dataset of safe positions $\mathcal{S}_{q}:=\{q_{i}\}_{i=1}^{N_{s}}$, and a dataset of unsafe positions $\mathcal{U}_{q}:=\{q_{i}\}_{i=1}^{N_{u}}$. Critically, the LiDAR does not provide safe or unsafe $\theta$ values, nor does it provide control inputs $u$. By joining the unsafe dataset $\mathcal{U}_{q}$ with unsafe positions $\mathcal{Z}_{q}$ sampled from beyond the convex hull of $\mathcal{S}_{q}$, we can perform Hamilton-Jacobi reachability analysis [9] to obtain a value function $V(x)$ that, for each $q\in\mathcal{S}_{q}$, is positive for safe $\theta$’s for which there exists a control signal that avoids $\mathcal{U}_{q}\cup\mathcal{Z}_{q}$. For each $q\in\mathcal{S}_{q}$ the safe angles are $\{\theta:V(q^{1},q^{2},\theta)\geq 0\}$. For a discretized range of $\theta$’s, the result is a dataset $\mathcal{D}:=\{(x_{i},u_{i})\}_{i=1}^{N}$ of safe states and control inputs, where the gradient of the value function, $\nabla V(x)$, is used to obtain safe control inputs according to $\pi^{*}(x)=\arg\max_{u\in U}\langle\nabla V(x),f(x)+g(x)u\rangle$. Note that synthesizing $\mathcal{D}$ into a locally-valid CBF effectively allows us to interpolate the gridded solution to the HJB-PDE, and to combine information from multiple PDE solutions by composing their learned CBFs.
@@ III-B Composing Locally-Valid CBFs
In the previous subsection, we generated datasets of safe state-action pairs, which can be used to obtain locally-valid CBFs via the Learning-CBF QP. Consider the task of composing $k$ locally-valid CBFs $h_{1},\dots,h_{k}$. Clearly, since $h_{i}$ is only constrained to be valid on some region $D_{i}$ where its dataset was sampled from, $h_{i}$ should be negative for $\mathbb{R}^{d_{x}}\setminus D_{i}$. Thus, if each CBF $h_{i}$ is valid on $D_{i}$, and $h_{i}(x)<0\ \forall x\in\mathbb{R}^{d_{x}}\setminus D_{i}$, then $H(x):=\max_{i\in\{1,\ldots,k\}}h_{i}(x)$ is a valid nonsmooth CBF on all $\bigcup_{i\in\{1,\ldots,k\}}D_{i}$.
Let $\phi_{s}(x,Z):\mathbb{R}^{d_{x}}\times\mathbb{R}^{d_{x}\times J}\to\mathbb{R}^{J}$ denote a vector of compactly supported radial basis functions that decay to zero in radius $s$. Specifically, each component $[\phi_{s}(x,Z)]_{j}$ of $\phi_{s}(x,Z)$ is a Wendland’s function of the form ${\phi}_{s}(x,z_{j}):=\bar{\phi}(\|x-z_{j}\|_{2}/s)=(1-r_{j}/s)_{+}^{p}f(r_{j}/s)$ for $r_{j}:=\|x-z_{j}\|_{2}$, where $(\cdot)_{+}:=\max(0,\cdot)$ and $f(r_{j}/s)$ is polynomial in $r_{j}/s$ with order less than $p$ [6]. Note that $\bar{\phi}(r_{j}/s)=0\ \forall r_{i}\geq s$.
Assume that $D_{i}$ is a compact and contains all centers $z_{j}$ of $h_{i}$. We show that as long as all centers $z_{j}$ whose corresponding weight $\theta_{j}$ is positive are at least a distance of $s$ away from $\mathrm{bd}(D_{i})$, which can be guaranteed by sampling unsafe points to increase the size of $\mathcal{U}$ in Eq. 3c, then $h_{i}(x)<0$ on all $\mathbb{R}^{d_{x}}\setminus D_{i}$. It follows that $\bigcup_{i}S_{i}$ is forward invariant under control inputs generated by $H(x)$. Let $d(z,D)$ denote the smallest Euclidean distance from point $z$ to set $D$.
@@
Consider $k$ continuously differentiable CBFs $\{h_{i}(x)\}_{i=1}^{k}$ parameterized as $h_{i}(x):=\phi_{s}(x,Z^{(i)})^{T}\theta^{(i)}-b$ with weights $\theta^{(i)}$, centers $Z^{(i)}$, and $b>0$. Let each $h_{i}(x)$ be locally valid on a corresponding region from $\{D_{i}\}_{i=1}^{k}$. For all $z_{j}^{(i)}\in Z^{(i)}$, assume that $z_{j}^{(i)}\in D_{i}$ and that $d(z_{j}^{(i)},\mathrm{bd}(D_{i}))\geq s$ if $\theta_{j}^{(i)}>0$. If for all $x\in\bigcup_{i\in\{1,\ldots,k\}}D_{i}$ with $H(x)\geq 0$ there exists $\epsilon>0$ such that the following (modified) CBF-QP has a bounded solution
$u^{*}(t)=\operatorname*{arg\,min}_{u\in U}$ $\quad\|u-u_{r}(t)\|_{2}^{2}$
$\mathrm{s.t.}$ $\quad\left\langle\nabla h_{i}(x),f(x)+g(x)u\right\rangle\geq-\alpha(H(x))$
$\mathllap{\forall i\in\{i^{\prime}:|h_{i^{\prime}}(x)-H(x)|\leq\epsilon\}},$
then $H(x)$ is a valid non-smooth CBF that guarantees forward invariance of the set $\bigcup_{i\in\{1,\ldots,k\}}S_{i}\subseteq\bigcup_{i\in\{1,\ldots,k\}}D_{i}$.
Proof: If $d(z_{j}^{(i)},\mathrm{bd}(D_{i}))\geq s$ for all $z_{j}^{(i)}$ s.t. $\theta_{j}^{(i)}>0$, then $b>0\Rightarrow h_{i}(x)<0$ (strict ineq.) $\forall x\in\mathbb{R}^{d_{x}}\setminus D_{i}$ by the compact support property of $\bar{\phi}_{s}$. By Theorem 3 of [5], it immediately follows that $H(x)$ is a non-smooth CBF that guarantees forward invariance of the set $\bigcup_{i\in\{1,\ldots,k\}}S_{i}$. ∎
The set $I_{\epsilon}(x):=\{i^{\prime}:|h_{i^{\prime}}(x)-H(x)|\leq\epsilon\}$ is called the “almost-active” set in [5]. The CBF-QP in Theorem 1 hence imposes a derivative constraint for each $h_{i}(x)$ with $i\in I_{\epsilon}(x)$ to obtain forward invariant control inputs.
Figure 4: Left: Local Wendland’s function CBF (green and blue contours) constrains the Dubins car system (magenta trajectory) to the safe region. A signed distance function cannot certify forward invariance for this system (gold trajectory, leaves safe set), even with unbounded actuation.
@@ IV Case Studies
To illustrate the validity of our approach, we incrementally learn global CBFs for two systems: a planar, feedback-linearizable system and a fixed-velocity Dubins car. For both case studies, we choose $s=1$ and obtain CS-RBFs $\phi_{s=1}(x,z_{j}):=\max\{0,1-r_{j}\}^{4}(1+4r_{j})/20$. For simplicity, we pick the centers $z_{j}$ to fall on a grid on their domains; more sophisticated arrangements of centers could be chosen, e.g. along safe trajectories provided by the oracle. The code for our case studies can be found at https://github.com/paullutkus/learning-cbfs-online.
IV-A Dubins Car
The dynamics of the Dubins car are $\dot{x}=(\dot{q}^{1},\dot{q}^{2},\dot{\theta})=(V\cos\theta,V\sin\theta,u)$ with fixed velocity $V:=0.1$ and initial condition $x(0):=(-1.1,-1.1,0)$. We also consider the input constraints $|u|\leq u_{\textrm{max}}:=0.4$. The environment consists of a central circular obstacle and four walls (Fig. 2).
For our idealized LiDAR sensor, we define a circular scan radius of $1.1$ (depicted in red in Fig. 4) around the position $q$ of the system. Because all states outside of the scan radius could potentially be obstacles, the incrementally learned CBF must keep the system within the scan radius while avoiding detected obstacles. Our local CBFs $h_{i}(x)$ and thus also the incremental CBF $H(x)$ accomplish this requirement, while satisfying input constraints, see Figs. 2, LABEL:, and 3.
@@ IV-A Dubins Car
In contrast, candidate CBFs constructed from sensor scans without incorporating dynamics may not be valid and fail to satisfy safety constraints. To illustrate this point, we consider signed distance functions (SDFs), which have been used as candidate CBFs in the past, see e.g.  [7]. Fig. 4 shows that an SDF cannot provide safety for this system, even with unbounded actuation ($u_{\text{max}}=\infty)$, because the gradient of an SDF to obstacles $l(x)$ is always orthogonal to the vector field of the Dubins car (since $\partial l/\partial\theta=0$). Though high-order CBFs can be used to resolve this incompatibility, we are able to control the system with a simpler, first-order CBF.
@@
Figure 5: Let $h_{i}|X_{i}$ denote that the CBF $h_{i}$ was obtained from data $X_{i}$. Top: The magnitudes of gradients of $\max_{i}\{h_{i}|X_{i}\}$ are visibly larger than those of $h|\bigcup_{i}X_{i}$, and that the zero-level set of $\max_{i}\{h_{i}|X_{i}\}$ is closer to the central obstacle. Bottom: The trajectory under $\max_{i}\{h_{i}|X_{i}\}$ reaches the target in less time, displaying larger actuation and speed while still maintaining safety. These results suggest that $\max_{i}\{h_{i}|X_{i}\}$ enjoys decreased conservatism.
IV-B Planar System
The dynamics of the planar system are:
$\dot{x}=\begin{bmatrix}1&0\\ 0&1\end{bmatrix}x+\begin{bmatrix}(x^{1})^{2}+\delta&0\\ 0&(x^{2})^{2}+\delta\end{bmatrix}u$
with $\delta:=0.33$ and initial condition $x(0)=(0,0)$. We consider the input constraints $|u|\leq u_{\textrm{max}}=1$. The environment consists of two vertically-aligned circular obstacles and four walls (Fig. 5). The scan radius of the LiDAR sensor is $1$.
@@ IV-B Planar System
Here, we compare our incremental CBF $H(x)$ with a CBF that is trained in a single-shot on all the data that was used to incrementally train $H(x)$. We find that not only do we match the performance of such a single-shot CBF, but our incremental CBF also performs better in this case (Fig. 5). Specifically, the zero-level set is closer to the obstacle, and larger gradients allow the system to approach a target more quickly. This is not surprising, as the maximum of continuously-differentiable CBFs is a more expressive function class than a single continuously differentiable CBF. Furthermore, incrementally learning a CBF is almost three times faster ($5.2$s total for four local CBFs vs. $14.9$s).

### v2

@@ III-A Obtaining Expert Demonstrations Online
Figure 2: Left-to-Right: Incremental learning of global Dubins car CBF, while safely exploring unknown environment under input constraints. LiDAR scans are taken where the trajectory changes color. Showcasing the correctness of local CBFs, each trajectory is constrained by the last-learned local CBF, whose zero-level set is shaded. At each $(q^{1},q^{2})$, red arcs denote unsafe angles $h(q^{1},q^{2},\theta)<0$. Blue arcs denote $h(q^{1},q^{2},\theta)\geq 0$.
In order to apply the Learning-CBF QP described in Section II-A we need access to a dataset of expert demonstrations containing safe state-action pairs (cf. Definition 2). This may not be possible during exploration. Suppose our measurement map is implemented via a LiDAR sensor. While LiDAR provides us with a point cloud of obstacles, it does not provide expert controls that avoid the obstacles. We provide two remedies for the lack of safe inputs. Our first solution is based on the dual norm. Suppose that the input set $U$ is a norm ball of the form $\|u\|\leq u_{\max}$. Then, the CBF derivative constraint $\sup_{u\in U}\langle\nabla h(x),\ f(x)+g(x)u\rangle\geq-\alpha(h(x))$ and hence the constraint Eq. 3 in the Learning-CBF QP can be rewritten in terms of the dual norm $\|u\|^{\star}=\sup_{x\in\mathbb{R}^{d_{n}}}\{\langle x,u\rangle:\|x\|\leq 1\}$ as follows [25]: $\langle\nabla h(x),f(x)\rangle+u_{\max}\|g(x)^{T}\nabla h(x)\|^{\star}+\alpha(h(x))\geq\gamma_{\text{dyn}}$. The above equation does not require safe inputs $u$, but results in a nonconvex Learning-CBF QP. A deeper problem is that sensors may not fully characterize the safe state (i.e. $d_{q}<d_{x}$). LiDAR on a car may provide safe & unsafe positions, but not safe steering angles and velocities, which are critical to avoid states that inevitably lead to collision. To address the limitations of the dual norm, our second approach defines the notion of a safety oracle, which takes the safe set returned by a sensor as input and returns safe state-action pairs for which there exists a control signal that keeps the system in this safe set for all time.
@@
Some examples of safety oracles are Hamilton-Jacobi reachability analysis [7], safe MPC [26], and obstacle-aware shooting methods [27]. Our proposed approach allows us to efficiently use local oracle computations in two ways. First, while oracle computations may only provide safe state-action pairs for a finite set of states, the learned locally-valid CBF interpolates between the given datapoints (cf. Definition 2), yielding safety guarantees over the entire local domain. Secondly, oracle computatations may be costly. If we recomputed the oracle over all previously seen states, the computational cost would grow in an unbounded manner as the agent explores. By computing the oracle over a local region around the agent, the computational cost stays constant during exploration. By synthesizing local safety computations into composable CBFs, we store and combine them in a principled way, circumventing the need to recalculate the safety of already-certified trajectories. Furthermore, our approach produces a safety filter (cf. Lemma 1) that can be used alongside any nominal controller, unlike many oracles alone. The following example illustrates our approach.
Figure 3: Global CBF for Dubins car. Pink trajectory of system under global CBF interpolates system behavior under local CBFs shown in Fig. 2.
Dubins car: Consider dynamics $\dot{x}=(\dot{q}^{1},\dot{q}^{2},\dot{\theta})=(v\cos\theta,v\sin\theta,u)$, with position $q:=(q^{1},q^{2})$ and fixed velocity $v$. The system operates in an unknown environment with access to a LiDAR sensor that returns safe and unsafe positions $\mathcal{S}_{q}:=\{q_{i}\}_{i=1}^{N_{s}}$ and $\mathcal{U}_{q}:=\{q_{i}\}_{i=1}^{N_{u}}$. Critically, the LiDAR does not provide safe or unsafe $\theta$ values, nor safe controls $u$. By joining the unsafe dataset $\mathcal{U}_{q}$ with unsafe positions $\mathcal{Z}_{q}$ sampled from beyond the convex hull of $\mathcal{S}_{q}$, we can perform Hamilton-Jacobi reachability analysis [7] to obtain a value function $V(x)$ that, for each $q\in\mathcal{S}_{q}$, is positive for safe $\theta$’s where there exists a control signal that avoids $\mathcal{U}_{q}\cup\mathcal{Z}_{q}$. For each $q\in\mathcal{S}_{q}$ the safe angles are $\{\theta:V(q^{1},q^{2},\theta)\geq 0\}$. For a discretized range of $\theta$’s, the result is a dataset $\mathcal{D}:=\{(x_{i},u_{i})\}_{i=1}^{N}$ of safe states and controls, where the gradient of the value function, $\nabla V(x)$, is used to obtain safe controls according to $\pi^{*}(x)=\arg\max_{u\in U}\langle\nabla V(x),f(x)+g(x)u\rangle$. Synthesizing $\mathcal{D}$ into a locally-valid CBF effectively interpolates the gridded solution to the HJB-PDE, and allows us to combine information from multiple solutions by composing their learned CBFs.
@@ III-B Composing Locally-Valid CBFs
In the previous subsection, we generated datasets of safe state-action pairs for obtaining locally-valid CBFs via the Learning-CBF QP. Consider the task of composing $k$ locally-valid CBFs $h_{1},\dots,h_{k}$. Because $h_{i}$ is only valid on $D_{i}$ where its dataset was sampled from, $h_{i}$ should be negative for $\mathbb{R}^{d_{x}}\setminus D_{i}$ in order to be globally well-defined. We show that if $h_{i}(x)<0\ \forall x\in\mathbb{R}^{d_{x}}\setminus D_{i}$, then $H(x):=\max_{i\in\{1,\ldots,k\}}h_{i}(x)$ is a valid nonsmooth CBF on all $\bigcup_{i\in\{1,\ldots,k\}}D_{i}$.
Let $\phi_{s}(x,Z):\mathbb{R}^{d_{x}}\times\mathbb{R}^{d_{x}\times J}\to\mathbb{R}^{J}$ be a vector of compactly-supported RBFs that decay to zero in radius $s$. Let each component $[\phi_{s}(x,Z)]_{j}$ of $\phi_{s}(x,Z)$ be a Wendland’s function of the form ${\phi}_{s}(x,z_{j}):=\bar{\phi}(\|x-z_{j}\|_{2}/s)=(1-r_{j}/s)_{+}^{p}f(r_{j}/s)$ for $r_{j}:=\|x-z_{j}\|_{2}$, where $(\cdot)_{+}:=\max(0,\cdot)$ and $f(r_{j}/s)$ is polynomial in $r_{j}/s$ with order less than $p$ [6]. Note that $\bar{\phi}(r_{j}/s)=0\ \forall r_{i}\geq s$. Now assume that $D_{i}$ is compact and contains all centers $z_{j}$ of $h_{i}$. We show that if all centers $z_{j}$ whose weight $\theta_{j}$ is positive are at least a distance of $s$ away from $\mathrm{bd}(D_{i})$, which can be guaranteed by sampling unsafe points to increase the size of $\mathcal{U}$ in Eq. 3c, then $h_{i}(x)<0$ on all $\mathbb{R}^{d_{x}}\setminus D_{i}$. It follows that $\bigcup_{i}S_{i}$ is forward invariant under controls generated by $H(x)$. The next assumptions precede our main result where $d(z,D)$ denotes the minimum $L_{2}$ distance from point $z$ to set $D$.
Assumption 1 (Persistent Feasibility)
$\exists\varepsilon>0$ such that the optimization problem $u^{*}(t)=\operatorname*{arg\,min}_{u\in U}\|u-u_{r}(t)\|_{2}^{2}$
subject $\text{subject to}\quad\left\langle\nabla h_{i}(x),f(x)+g(x)u\right\rangle\geq-\alpha(H(x))$
$\mathllap{\forall i\in\{i^{\prime}:|h_{i^{\prime}}(x)-H(x)|\leq\epsilon\}},$ (4)
has a bounded solution for all $x\in\bigcup_{i\in\{1,\ldots,k\}}S_{i}$.
Even if Eq. 4 is infeasible, safety is never jeopardized, since the agent can simply stay in the current CBF, and reattempt crossing into a neighboring CBF at a later time. Note the assumption of persistent feasibility is also made in [5].
Assumption 2 (RBF Decay Buffer)
The centers $Z^{(i)}$ of CBF $h_{i}$ satisfy the following: For all $z_{j}^{(i)}\in Z^{(i)}$, $z_{j}^{(i)}\in D_{i}$ and $d(z_{j}^{(i)},\mathrm{bd}(D_{i}))\geq s$ if $\theta_{j}^{(i)}>0$.
This assumption is easily satisfied by adding negativity constraints Eq. 3c for points sampled in a spherical shell of width $s$, with inner radius equal to the outermost center. This a sufficient but not necessary condition, and we find that a shell of width $s^{\prime}\ll s$ is sufficient in our experiments.
@@
Consider $k$ continuously differentiable CBFs $\{h_{i}(x)\}_{i=1}^{k}$ parameterized as $h_{i}(x):=\phi_{s}(x,Z^{(i)})^{T}\theta^{(i)}-b$ with weights $\theta^{(i)}$, centers $Z^{(i)}$, and $b>0$. Let each $h_{i}(x)$ be locally-valid on a corresponding region from $\{D_{i}\}_{i=1}^{k}$. If 1 and 2 hold, then $H(x)=\max_{i}\{h_{i}(x)\}$ is a valid non-smooth CBF that guarantees forward invariance of the set $\bigcup_{i\in\{1,\ldots,k\}}S_{i}\subseteq\bigcup_{i\in\{1,\ldots,k\}}D_{i}$.
Proof: If $d(z_{j}^{(i)},\mathrm{bd}(D_{i}))\geq s$ for all $z_{j}^{(i)}$ s.t. $\theta_{j}^{(i)}>0$, then $b>0\Rightarrow h_{i}(x)<0$ (strict ineq.) $\forall x\in\mathbb{R}^{d_{x}}\setminus D_{i}$ by the compact support property of $\bar{\phi}_{s}$. By [5] Thm. 3 and convexity of the almost-active gradient, $H(x)$ is a non-smooth CBF yielding forward invariance for the set $\bigcup_{i\in\{1,\ldots,k\}}S_{i}$. ∎
The set $I_{\epsilon}(x):=\{i^{\prime}:|h_{i^{\prime}}(x)-H(x)|\leq\epsilon\}$ is called the “almost-active” set in [5]. The QP in Theorem 1 thus imposes a derivative constraint for each $h_{i}(x)$ with $i\in I_{\epsilon}(x)$.
Figure 4: CS-RBF barrier (teal contours) constrains Dubins car (pink) to the safe region. Signed distance functions cannot certify fwd. invariance for this system (gold, leaves safe set), even with unbounded actuation.
@@ IV Case Studies
To illustrate the validity of our approach, we incrementally learn global CBFs for two systems: a planar, feedback-linearizable system and a fixed-velocity Dubins car. For both case studies, we choose $s=1$ and obtain CS-RBFs $\phi_{s=1}(x,z_{j}):=\max\{0,1-r_{j}\}^{4}(1+4r_{j})/20$. For simplicity, we pick the centers $z_{j}$ to fall on a grid in $D_{i}$; more sophisticated arrangements of centers could be chosen, e.g. along safe trajectories provided by the oracle. The code for our case studies can be found at https://github.com/paullutkus/learning-cbfs-online.
@@
Dubins Car: The dynamics of the Dubins car are $\dot{x}=(\dot{q}^{1},\dot{q}^{2},\dot{\theta})=(v\cos\theta,v\sin\theta,u)$ with fixed velocity $v:=0.1$, initial condition $x(0):=(-1.1,-1.1,0)$, and input constraints $|u|\leq u_{\textrm{max}}:=0.4$. The environment consists of a central obstacle and four walls (Fig. 2). For our idealized LiDAR, we define a circular scan radius of $1.1$ (depicted in red in Fig. 4) centered on the agent. Since all states outside of the scan could be obstacles, the local CBF must keep the system within this radius. Our local CBFs $h_{i}(x)$ and composite CBF $H(x)$ yield invariance while satisfying input constraints (see Figs. 2, LABEL:, and 3). In contrast, candidate CBFs constructed without incorporating dynamics may not be valid nor provide safety. To illustrate this, we consider signed distance functions (SDFs), which have been used as candidate CBFs in the past, see e.g. [19]. Fig. 4 shows that an SDF cannot provide safety for this system, even with unbounded actuation, as the gradient of an SDF to obstacles $l(x)$ is always orthogonal to the vector field of the Dubins car ($\partial l/\partial\theta=0$). While high-order CBFs can resolve this incompatibility, we are able to control the system with a simpler, first-order CBF. We also perform a timing comparison with [10], assessing the cost of local CBFs when using an HJB oracle. We find a $45\%$ average overhead compared to using local HJB only (2.45s vs 3.55s), suggesting that a composite CBF can be obtained for minimal added cost.
@@
Figure 5: $h_{i}|X_{i}$ denotes that $h_{i}$ was obtained from data $X_{i}$. Top: The gradient norms of $\max_{i}\{h_{i}|X_{i}\}$ are visibly larger than those of $h|\bigcup_{i}X_{i}$, and the zero-level set of $\max_{i}\{h_{i}|X_{i}\}$ is closer to the central obstacle. Bottom: Trajectory under $\max_{i}\{h_{i}|X_{i}\}$ reaches target faster, with larger actuation & speed while maintaining safety, suggesting decreased conservatism.
@@
Planar System: The dynamics of the planar system are $(\dot{x}_{1},\dot{x}_{2})=(x_{1}+(x_{1}^{2}+\delta)u_{1},x_{2}+(x_{2}^{2}+\delta)u_{2})$ with $\delta:=0.33$ and initial condition $x(0)=(0,0)$. We consider the input constraints $|u|\leq u_{\textrm{max}}=1$. The environment consists of two vertically-aligned circular obstacles and four walls (Fig. 5). The scan radius of the LiDAR sensor is $1$. Here, we compare our incremental CBF $H(x)$ with a CBF trained in a single-shot on all the data that was used to incrementally train $H(x)$. Not only do we match performance with such a single-shot CBF, but our incremental CBF exceeds it in this case (Fig. 5). Specifically, the zero-level set is closer to the obstacle, and larger gradients allow the system to approach a target faster. This is not surprising, as the max of continuously-differentiable CBFs is a more expressive function class than single continuously-differentiable CBFs. Furthermore, incremental learning is almost three times faster ($5.2$s total for four local CBFs vs. $14.9$s for one global).